### v2

Abstract.
We investigate the class of $\mathcal{K}$}-Lorentzian polynomials, which generalizes the well-known class of Lorentzian polynomials. As demonstrated in [GPlorentzian], these polynomials are equivalent to $\mathcal{K}$}-completely log-concave ($\mathcal{K}$}-CLC) forms. Exploiting the structure of directional derivatives and Hessians, we prove that the Hessian matrices of $\mathcal{K}$}-CLC forms satisfy a generalized Perron–Frobenius property over a self-dual cone $\mathcal{K}$}: they admit a distinguished eigenvector in $\operatorname{int}{\mathcal{K}}$. We introduce a line-restriction perspective on $\mathcal{K}$}-CLC. This leads to a hierarchy of log-concavity and Newton-type log-concavity inequalities for the coefficients of univariate restrictions. This provides useful necessary conditions for $\mathcal{K}$}-CLC in terms of directional derivatives.
As a key application, we derive degree-dependent criteria for Hurwitz stability of $\mathcal{K}$}-CLC univariate restrictions. In particular, we prove that all strictly $\mathcal{K}$}-Lorentzian polynomials of degree $d\leq 4$ are Hurwitz-stable over $\mathcal{K}$}, and we obtain refined inequalities that give sufficient conditions for Hurwitz stability in degrees $d\geq 5$. These results recover and extend known stability criteria in the classical Lorentzian case ${\mathcal{K}}={\mathbb{R}}^{n}_{+}$ to the general conic setting. Moreover, these Hurwitz–stability and Newton-type log-concavity coefficient criteria yield simple sufficient conditions for asymptotic stability of linear time-invariant systems via their characteristic polynomials.
Key words and phrases:
$\mathcal{K}$}-Lorentzian polynomials, $\mathcal{K}$}-completely log-concave polynomials, convex cones, log-concavity, Newton-type inequalities, Perron–Frobenius theorem, Hurwitz stability, linear time-invariant (LTI) systems
2020 Mathematics Subject Classification:
Primary 15A18; Secondary 26B25, 52A20, 34D20, 93D05, 15B48.
1. Introduction
$\mathcal{K}$}-Lorentzian polynomials were introduced and studied by Brändén and Leake in [Leake], and by Blekherman and Dey in [GPlorentzian]. The approach in [GPlorentzian] not only generalizes Lorentzian polynomials introduced by Huh and Brändén [Brandenlorentzian] and, independently, by Anari, Oveis Gharan and Vinzant under the name of completely log-concave polynomials [Cynthialog], but also provides a unified framework for extending this class to arbitrary proper convex cones.
In this work we use the term $\mathcal{K}$}-Lorentzian for the homogeneous setting and $\mathcal{K}$}-completely log-concave ($\mathcal{K}$}-CLC) for the possibly non-homogeneous setting. A central result of [GPlorentzian] is that, for any proper cone $\mathcal{K}$}, the class of $\mathcal{K}$}-Lorentzian forms coincides with the class of $\mathcal{K}$}-CLC forms of the same degree; see Theorem˜2.9. This gives a powerful bridge between convex-analytic log-concavity on cones and the algebraic signature conditions that define Lorentzian behavior via iterated directional derivatives and quadratic forms.
Our first goal in this part is to refine the structural picture of $\mathcal{K}$}-Lorentzian and $\mathcal{K}$}-CLC polynomials. For homogeneous forms, we show that the Hessian matrices evaluated on $\mathcal{K}$} have a very rigid eigenvalue structure: they have exactly one positive eigenvalue, and, over self-dual cones, satisfy a generalized Perron–Frobenius property (Theorem˜3.6, Theorem˜3.7, Corollary˜3.8). This connects $\mathcal{K}$}-Lorentzian polynomials to $\mathcal{K}$}-nonnegative and $\mathcal{K}$}-irreducible matrices, and to the cone-theoretic Perron–Frobenius theory developed in [Kreinlinear, Barkeralgebraic].
Next, we introduce a linewise notion of $\mathcal{K}$}-complete log-concavity: for each $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}$, we consider the univariate restriction
$t\longmapsto f(x+tv),$
and require all iterated derivatives in $t$ to be log-concave on the appropriate interval. We show that $\mathcal{K}$}-CLC in the sense of Definition˜2.1 implies linewise $\mathcal{K}$}-CLC (Proposition˜3.13). Combined with sharp univariate characterizations of complete log-concavity, this yields log-concavity of the factorial-weighted coefficient sequences of $f(x+tv)$ and, consequently, log-concavity of the directional derivative sequence
$\bigl\{f(x),D_{v}f(x),D_{v}^{2}f(x),\dots,D_{v}^{d}f(x)\bigr\}$
for any $x,v\in\operatorname{int}{\mathcal{K}}$ (Theorem˜3.16). We also obtain positivity of the coefficients of these univariate restrictions for strictly $\mathcal{K}$}-CLC polynomials (Proposition˜3.14, Corollary˜3.15).
Then we connect these structural inequalities to classical Hurwitz stability of univariate polynomials. Using Routh–Hurwitz and Liénard–Chipart criteria together with the log-concavity relations among coefficients, we show that strict $\mathcal{K}$}-CLC implies Hurwitz stability in low degrees. Specifically, we prove that every univariate strictly CLC polynomial of degree $d\leq 4$ with positive coefficients is Hurwitz-stable (Proposition˜4.14), and we obtain refined sufficient conditions for Hurwitz stability in degree $d=5$ (Theorem˜4.17) based on strengthened log-concavity inequalities.
In the final part of this paper our Newton-type coefficient inequalities for strictly CLC and strictly $\mathcal{K}$}-Lorentzian polynomials yield simple sufficient tests ensuring stability of a linear time-invariant system (LTI) directly from the coefficients of $\chi_{A}(t)$, (Theorem˜5.2 and Theorem˜5.4).
Organization. The rest of this paper is structured as follows. In Section˜2 we recall basic definitions on proper cones, $\mathcal{K}$}-nonnegativity, $\mathcal{K}$}-irreducibility, and the equivalence between $\mathcal{K}$}-Lorentzian and $\mathcal{K}$}-CLC forms. In Section˜3 we develop several key properties of $\mathcal{K}$}-Lorentzian polynomials, including Hessian characterizations, cone-theoretic Perron–Frobenius conditions, and a line-restriction framework for $\mathcal{K}$}-CLC, leading to log-concavity of directional derivatives. In Section˜4 we relate these inequalities to the classical half-plane property and Hurwitz stability for univariate restrictions, deriving degree-dependent criteria for stability. Section˜5 applies the Hurwitz–stability criteria and Newton-type log-concavity inequalities developed earlier to the characteristic polynomials of linear time-invariant (LTI) systems, yielding simple sufficient conditions for (asymptotic) stability expressed directly in terms of their coefficients.
2. State of the Art
A nonempty convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{n}$ is said to be a cone if $c{\mathcal{K}}\subseteq{\mathcal{K}}$ for all $c\geq 0$. A cone $\mathcal{K}$} is called proper if it is closed (in the Euclidean topology on ${\mathbb{R}}^{n}$), pointed (i.e., ${\mathcal{K}}\cap(-{\mathcal{K}})=\{0\}$), and solid (i.e., the topological interior of $\mathcal{K}$}, denoted as $\operatorname{int}{\mathcal{K}}$, is nonempty).
Let ${\mathbb{R}}[x]$ represent the space of $n$-variate polynomials over $\mathbb{R}$}, and ${\mathbb{R}}[x]_{\leq d}$ represent the space of $n$-variate polynomials over $\mathbb{R}$} with degree at most $d$ and ${\mathbb{R}}[x]^{d}_{n}$ denote the set of real homogeneous polynomials (aka forms) in $n$ variables of degree $d$. A polynomial $f\in{\mathbb{R}}[x]$ is said to be (strictly) log-concave at a point $a\in{\mathbb{R}}^{n}$ if $f(a)>0$, and $\log f$ is a (strictly) concave function at $a$, i.e. the Hessian of $\log f$ is negative semidefinite (negative definite) at $a$. A polynomial $f\in{\mathbb{R}}[x]$ is log-concave on a proper cone ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ if $f$ is log-concave at every point of the interior of $\mathcal{K}$}. By convention, the zero polynomial is log-concave (but not strictly log-concave) at all points of ${\mathbb{R}}^{n}$.
If ${\mathcal{K}}={\mathcal{K}}^{\ast}$, a self-dual cone, then $K$ is closed, pointed, and full dimensional[Barkerselfdual]. We denote quadratic forms by a lower case letter, and the matrix of the quadratic form by the corresponding upper case letter, i.e. if $q$ is a quadratic form, then its matrix is $Q$ and $q(x)=x^{t}Qx$.
For a point $a\in{\mathbb{R}}^{n}$ and $f\in{\mathbb{R}}[x]$, $D_{a}f$ denotes the directional derivative of $f$ in direction $a$: $D_{a}f=\sum_{i=1}^{n}a_{i}\frac{\partial f}{\partial x_{i}}$. Here are the definitions of the $\mathcal{K}$}-completely log-concave polynomials and $\mathcal{K}$}-Lorentzian forms.
Definition 2.1.
[GPlorentzian] A polynomial (form) $f\in{\mathbb{R}}[x]_{\leq d}$ is called a $\mathcal{K}$}-completely log-concave aka $\mathcal{K}$}-CLC (form) on a proper convex cone $\mathcal{K}$} if for any choice of $a_{1},\dots,a_{m}\in{\mathcal{K}}$, with $m\leq d$, we have that $D_{a_{1}}\dots D_{a_{m}}f$ is log-concave on $\operatorname{int}{\mathcal{K}}$. A polynomial (form) $f\in{\mathbb{R}}[x]$ is strictly $\mathcal{K}$}-CLC if for any choice of $a_{1},\dots,a_{m}\in{\mathcal{K}}$, with $m\leq d$, $D_{a_{1}}\dots D_{a_{m}}f$ is strictly log-concave on all points of $\mathcal{K}$}.
Definition 2.2.
[GPlorentzian] Let $\mathcal{K}$} be a proper convex cone. A form $f\in{\mathbb{R}}[x]_{n}^{d}$ of degree $d\geq 2$ is said to be $\mathcal{K}$}-Lorentzian if for any $a_{1},\dots,a_{d-2}\in\operatorname{int}{\mathcal{K}}$, the quadratic form $q=D_{a_{1}}\dots D_{a_{d-2}}f$ satisfies the following conditions:
  1. (1)
    The matrix $Q$ of $q$ has exactly one positive eigenvalue.
  2. (2)
    For any $x,y\in\operatorname{int}{\mathcal{K}}$ we have $y^{t}Qx=\langle y,Qx\rangle>0$.
For degree $d\leq 1$ a form is $\mathcal{K}$}-Lorentzian if it is nonnegative on $\mathcal{K}$}.
Definition 2.3.
Let $\mathcal{K}$} be a proper convex cone in ${\mathbb{R}}^{n}$. An $n\times n$ matrix $A$ is $\mathcal{K}$}-nonnegative if $Ax\in{\mathcal{K}}$ for any $x\in{\mathcal{K}}$, i.e. $A({\mathcal{K}})\subseteq{\mathcal{K}}$. A matrix $A$ is $\mathcal{K}$}-positive if $Ax\in\operatorname{int}{\mathcal{K}}$ for any nonzero $x\in{\mathcal{K}}$. A $\mathcal{K}$}-nonnegative matrix $A$ is called $\mathcal{K}$}-irreducible if $A$ leaves no (proper) face of $\mathcal{K}$} invariant. A $\mathcal{K}$}-nonnegative matrix which is not $\mathcal{K}$}-irreducible is called $\mathcal{K}$}-reducible, cf. [gantmacher], [Vandergraftspectral].
Lemma 2.4.
[GPlorentzian, Lemma 3.1] A quadratic form $q(x)=x^{t}Qx\in{\mathbb{R}}[x]_{n}^{2}$ is $\mathcal{K}$}-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and one of the following conditions holds:
  1. (a)
    for all $x\in{\mathcal{K}}$ we have $Qx\in{\mathcal{K}}^{*}$, i.e. $Q({\mathcal{K}})\subseteq{{\mathcal{K}}^{*}}$.
  2. (b)
    $y^{t}Qx\geq 0$ for all $x,y$ spanning extreme rays of $\mathcal{K}$}.
  3. (c)
    $x^{t}Qx>0$ for all $x\in\operatorname{int}{\mathcal{K}}$.
Similar to copositive matrices over the nonnegative orthant, we define $\mathcal{K}$}-copositive matrices over $\mathcal{K}$}, see [stabilitylevi]. A matrix $Q$ is (strictly) $\mathcal{K}$}-copositive if $x^{t}Qx(>0)\geq 0$ for all $x\in{\mathcal{K}}$.
Corollary 2.5.
If a quadratic form $q(x)=x^{t}Qx\in{\mathbb{R}}[x]_{n}^{2}$ is (strictly) $\mathcal{K}$}-Lorentzian, then $Q$ is (strictly) $\mathcal{K}$}-copositive matrix.
For a self-dual cone $\mathcal{K}$} we have the following result.
Corollary 2.6.
[GPlorentzian, Cor 3.8] Let $\mathcal{K}$} be a self-dual cone. Then the quadratic form $q=x^{t}Qx$ is $\mathcal{K}$}-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is $\mathcal{K}$}-nonnegative.
Proposition 2.7.
[GPlorentzian, Cor 3:10] Let ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a self-dual cone. Then $q=x^{t}Qx$ is $\mathcal{K}$}-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is either nonsingular and $\mathcal{K}$}-irreducible, or singular and $\mathcal{K}$}-nonnegative.
Remark 2.8.
Then $q$ is a quadratic Lorentzian (aka stable) polynomial if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is nonnegative matrix.
Furthermore, we use the following equivalence result between $\mathcal{K}$}-Lorentzian and $\mathcal{K}$}-CLC polynomials, proved in [GPlorentzian, Theorem 4:10].
Theorem 2.9.
Let ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a proper cone. A form $f\in{\mathbb{R}}[x]_{n}^{d}$ is $\mathcal{K}$}-Lorentzian if and only if $f$ is $\mathcal{K}$}-CLC. Equivalently, $\operatorname{\mathrm{SL}}^{d}({\mathcal{K}})=\mathrm{CLC}^{d}({\mathcal{K}})$.
3. Properties of $\mathcal{K}$}-Lorentzain polynomials
3.1. Proper Convex Cones
Using Lemma˜2.4 we show that $\mathcal{K}$}-nonnegativity is a sufficient condition for $\mathcal{K}$}-Lorentzian polynomials if $\mathcal{K}$} is contained in its dual cone, $\mathcal{K}$$^{\ast}$.
Corollary 3.1.
Consider the quadratic form $q(x)=x^{t}Qx\in{\mathbb{R}}[x]_{n}^{2}$ and $\mathcal{K}$} be a proper cone such that ${\mathcal{K}}\subseteq{\mathcal{K}}^{\ast}$. If the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is $\mathcal{K}$}-nonnegative, then $q$ is a $\mathcal{K}$}-Lorentzian on the proper convex cone $\mathcal{K}$}.
Proof.
Since $Q$ is $\mathcal{K}$}-nonnegative, so $Q({\mathcal{K}})\subseteq{\mathcal{K}}\subseteq{\mathcal{K}}^{\ast}$. Then the rest follows from Lemma˜2.4. ∎
Based on Lemma˜2.4 we show that $\mathcal{K}$}-nonnegativity is a necessary condition for $\mathcal{K}$}-Lorentzian polynomials if $\mathcal{K}$} contains its dual cone, $\mathcal{K}$$^{\ast}$.
Corollary 3.2.
If a quadratic form $q(x)\in{\mathbb{R}}[x]_{n}^{2}$ is $\mathcal{K}$}-Lorentzian on a proper convex cone $\mathcal{K}$} and ${\mathcal{K}}^{\ast}\subseteq{\mathcal{K}}$, the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is $\mathcal{K}$}-nonnegative.
Proof.
By Lemma˜2.4 we have the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q({\mathcal{K}})\subseteq{{\mathcal{K}}^{*}}$. Thus, $Q({\mathcal{K}})\subseteq{\mathcal{K}}^{*}\subseteq{\mathcal{K}}$. Therefore, $Q$ is $\mathcal{K}$}-nonnegative by Definition˜2.3. ∎
Remark 3.3.
The converse to Corollary˜3.2 is not true. It’s easy to construct counterexamples where ${\mathcal{K}}^{\ast}\subseteq{\mathcal{K}}$ and $Q({\mathcal{K}})\subseteq{\mathcal{K}}$ but $Q({\mathcal{K}})\not\subseteq{\mathcal{K}}^{\ast}$.
It’s shown that any closed convex cone contains its dual cone, i.e., ${\mathcal{K}}^{\ast}\subset{\mathcal{K}}$, if and only if there exist $y,z\in{\mathcal{K}}$ such that x=yzx=y-z and $\langle y,z\rangle=0$, cf. [Remarkcopositive] for detail. This is used to show that every $x\in{\mathbb{R}}^{n}$ has an (unique) orthogonal decomposition on $\mathcal{K}$} in [Cross] which have many well-known consequences. Additionally, the sum of two $\mathcal{K}$}-nonnegative matrices is $\mathcal{K}$} nonnegative. Furthermore, if $A$ and $B$ are $\mathcal{K}$}-nonnegative and $A$ is $\mathcal{K}$}-irreducible, then so is $A+B$, cf. [Bermannonnegative, Chap-1]. However, the sum of two log-concave polynomials may not be log-concave. The sufficient condition for the sum of two $\mathcal{K}$}-Lorentzian polynomials to also be $\mathcal{K}$}-Lorentzian can be found in [Cynthialog3, Lemma 3.3] for the nonnegative orthant, and [GPlorentzian, Theorem 4.15] for any proper convex cone $\mathcal{K}$}.
Theorem 3.4.
[GPlorentzian, Theorem 4.15] Let $f,g\in{\mathbb{R}}^{d}_{n}[x]$ be two $\mathcal{K}$}-Lorentzian polynomials over the self-dual cone $\mathcal{K}$}. Then the sum $f+g$ is $\mathcal{K}$}-Lorentzian if there exist vectors $b,c\in{\mathcal{K}}$ such that $D_{b}f=D_{c}g\not\equiv 0$.
3.2. Perron-Frobenius Theorem for Cones
Let $\mathcal{K}$} be a proper convex cone. We can leverage the relationship between $\mathcal{K}$}-Lorentzian polynomials and the generalized Perron-Frobenius theorem [Kreinlinear], [Barkeralgebraic] to distinguish potential $\mathcal{K}$}-Lorentzian polynomial candidates from given log-concave polynomials. This also provides valuable insight into the structure of the cone K.
Theorem 3.5.
(Generalization of Perron-Frobenius Theorem) [Kreinlinear], [Barkeralgebraic]: Let $A$ be a $\mathcal{K}$}-irreducible matrix with spectral radius $\rho$. Then
  1. (1)
    $\rho$ is a simple positive eigenvalue of $A$,
  2. (2)
    There exists a (up to a scalar multiple) unique $\mathcal{K}$}-positive (right) eigenvector $u$ of $A$ corresponding to $\rho$,
  3. (3)
    $u$ is the only $\mathcal{K}$}-semipositive eigenvector for $A$ (for any eigenvalue),
  4. (4)
    ${\mathcal{K}}\cap(\rho I-A){\mathbb{R}}^{n}=\{0\}$.
By applying [GPlorentzian, Lemma 4.12] and [GPlorentzian, Prop:4.13] we derive the following characterization of the Hessian matrices evaluated at any point over the cone $\mathcal{K}$} for any $\mathcal{K}$}-Lorentzian polynomials where $\mathcal{K}$} is a proper convex cone.
Theorem 3.6.
Let $f\in{\mathbb{R}}[x]_{n}^{d}$ be a nonzero form of degree $d\geq 2$, and $\mathcal{K}$} be a proper convex cone in ${\mathbb{R}}^{n}$.
  1. (a)
    If $f$ is a strictly $\mathcal{K}$}-CLC form, then $H_{f}(a)$ is nonsingular and has exactly one positive eigenvalue for all $a\in{\mathcal{K}}$. Also the quadratic form $x^{t}H_{f}(a)x$ is negative definite on $(H_{f}(a)b)^{\perp}$ for every $a,b\in{\mathcal{K}}$ such that $H_{f}(a)b\neq 0$.
  2. (b)
    If $f$ is a $\mathcal{K}$}-CLC form, then $H_{f}(a)$ has exactly one positive eigenvalue for all $a\in\operatorname{int}{\mathcal{K}}$. Also the quadratic form $x^{t}H_{f}(a)x$ is negative semidefinite on $(H_{f}(a)b)^{\perp}$ for every $a\in\operatorname{int}{\mathcal{K}}$, and $b\in{\mathcal{K}}$ such that $H_{f}(a)b\neq 0$.
In special cases using the Definition˜2.2 and [GPlorentzian, Them:5.1] we have some necessary conditions for $\mathcal{K}$}-Lorentzian polynomials over the self-dual cone $\mathcal{K}$}.
Theorem 3.7.
Let $f\in{\mathbb{R}}[x]_{n}^{d}$ be a $\mathcal{K}$}-Lorentzian polynomial of degree $d\geq 2$ over the self-dual cone $\mathcal{K}$}. Then for all $a\in{\mathcal{K}}$
  1. (a)
    $H_{f}(a)$ is either nonsingular and $\mathcal{K}$}-irreducible, or singular and $\mathcal{K}$}-nonnegative.
  2. (b)
    $H_{f}(a)$ has exactly one positive eigenvalue.
By Theorem˜3.5 we have more stronger necessary conditions for $\mathcal{K}$}-Lorentzian polynomials over the self-dual cone $\mathcal{K}$}.
Corollary 3.8.
Let $f\in{\mathbb{R}}^{d}_{n}[x]$ be a $\mathcal{K}$}-Lorentzian polynomial over the self-dual cone $\mathcal{K}$}. Then for any $a\in{\mathcal{K}}$, the nonsingular Hessian $H_{f}(a)$ satisfies generalized Perron-Frobenius property.
Proof.
Since $f$ is a $\mathcal{K}$}-Lorentzian polynomial, so for any $a\in{\mathcal{K}}$, the nonsingular Hessian $H_{f}(a)$ is $\mathcal{K}$}-irreducible matrix and has exactly one positive eigenvalue. Therefore, if $H_{f}(a)$ is nonsingular, and the positive eigenvalue corresponds to the spectral radius, then by Theorem˜3.5, its corresponding eigenvector must lie in the interior of the cone $\mathcal{K}$}. ∎
The geometry of the set of symmetric Perron–Frobenius matrices has been studied in [SymmetricPFstructure]. In fact, it is not convex but is star convex with the identity matrix as the center. It’s a natural question to describe the geometry of the set of symmetric Perron–Frobenius matrices which have exactly one positive eigenvalue. We leave this question for future work.
3.3. Properties of $\mathcal{K}$}-CLC polynomials
Here we propose another definition for $\mathcal{K}$}-CLC multivariate polynomials by restricting them on certain lines. This allows us to show the log-concavity property among the coefficients of the corresponding univariate restrictions of a $\mathcal{K}$}-CLC polynomial. Recall that a univariate polynomial $f(t)$ is log-concave over an interval $I$ if $f(t)>0$ for all $t\in I$ and $f(t)f^{{}^{\prime\prime}}(t)\leq(f^{{}^{\prime}}(t))^{2}$ for all $t\in I$. Equivalently, a univariate polynomial $f(t)$ is log-concave over an interval $I$ if $f(t)>0$ for all $t\in I$ and $\frac{d^{2}}{dt^{2}}\log f(t)\leq 0$ for all $t\in I$. So, we can define $\mathcal{K}$}-CLC for any real multivariate polynomial as in Definition˜2.1, need not be homogeneous. If they are homogeneous, the set of $\mathcal{K}$}-Lorentzian polynomials of degree $d$ are equivalent to the set of $\mathcal{K}$}-CLC forms , cf. [GPlorentzian]). On the other hand, over the nonnegative orthant, a characterization for bivariate $\mathcal{K}$}-CLC form has been established in the literature, providing a resolution to Mason’s conjecture.
Theorem 3.9.
[Gurvitsmultivariate], [Cynthialog3] $f=\sum_{k=0}^{n}c_{k}x^{n-k}y^{k}\in{\mathbb{R}}[x,y]$ is completely log- concave if and only if the sequence of nonnegative coefficients, $\{c_{0},\dots,c_{n}\}$ is ultra log-concave, i.e., for every $1<k<n$, $\left(\frac{c_{k}}{\binom{n}{k}}\right)^{2}\geq\frac{c_{k-1}}{\binom{n}{k-1}}\frac{c_{k+1}}{\binom{n}{k+1}}$.
Lemma 3.10.
Let $f(t)=\sum_{i=0}^{d}a_{i}t^{i}$ be a univariate polynomial. The univariate polynomial $f$ is CLC over $t\geq 0$ if and only if $a_{i}>0$ and the sequence $\{0!a_{0},\dots,d!a_{d}\}$ is a log-concave sequence, i.e., $ia_{i}^{2}\geq(i+1)a_{i-1}a_{i+1}$ for all $i=0(1)d$.
Proof.
It’s straightforward to verify that $f(t)=\sum_{i=0}^{d}a_{i}t^{i}$ with $a_{i}>0$ is CLC for $t\geq 0$ if and only if its homogenization $f(t,u)=\frac{u^{d}}{d!}a_{0}+\frac{u^{d-1}}{(d-1)!}a_{1}t+\dots+\frac{u^{0}}{0!}a_{d}t^{d}$ is CLC over ${\mathbb{R}}^{2}_{\geq 0}$. Then the rest follows from Theorem˜3.9 and the identity $a_{i}^{2}(i!)^{2}\geq a_{i-1}a_{i+1}(i-1)!(i+1)!\Leftrightarrow ia_{i}^{2}\geq(i+1)a_{i-1}a_{i+1}$. ∎
Corollary 3.11.
The polynomial $f(t)=\sum_{i=0}^{d}a_{i}t^{d-i}$ with $a_{i}>0$ is CLC for all $t\geq 0$ if and only $\{d!a_{0},\dots,i!a_{d-i},\dots,0!a_{d}\}$ is a log-concave sequence
Proof.
Due to Lemma˜3.10 and symmetry, $((d-k)!a_{k})^{2}\geq(d-k-1)!(d-k+1)!a_{k-1}a_{k+1}\Leftrightarrow(d-k)a_{k}^{2}\geq(d-k+1)a_{k-1}a_{k+1}\Leftrightarrow ia_{d-i}^{2}\geq(i+1)a_{d-i-1}a_{d-i+1}$. ∎
It is well known that convexity can be characterized in terms of one-dimensional convexity. More precisely, a function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is convex over a convex set $\mathcal{K}$} if and only if for any $x\in{\mathcal{K}}$ and $v\in{\mathbb{R}}^{n}$, the function $g:{\mathbb{R}}\rightarrow{\mathbb{R}},g(t)=f(x+tv)$ is convex as function in $t$ for all $t$ such that $x+tv\in{\mathcal{K}}$ [Boydconvex]. In the same spirit, we introduce a line-restriction notion for log-concavity and complete log-concavity over cones.
Definition 3.12 (Linewise $\mathcal{K}$}-completely log-concave).
Let ${\mathcal{K}}\subseteq{\mathbb{R}}^{n}$ be a proper convex cone and let $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$. For $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, set
$I_{x,v}\coloneqq\{\,t\in{\mathbb{R}}:x+tv\in\operatorname{int}{\mathcal{K}}\,\},\qquad g_{x,v}(t)\coloneqq f(x+tv),\quad t\in I_{x,v}.$
  • We say that $f$ is linewise log-concave over $\mathcal{K}$} if, for every $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, the univariate function $g_{x,v}$ is log-concave on $I_{x,v}$.
  • We say that $f$ is linewise $\mathcal{K}$}-completely log-concave (linewise $\mathcal{K}$}-CLC) if, for every $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, the univariate polynomial $g_{x,v}$ is completely log-concave on $I_{x,v}$ in the univariate sense (i.e., for every $m\geq 0$, the derivative $g_{x,v}^{(m)}$ is log-concave on $I_{x,v}$).
We next show that $\mathcal{K}$}-complete log-concavity in the sense of Definition˜2.1 implies the linewise property from Definition˜3.12.
Proposition 3.13.
Let ${\mathcal{K}}\subseteq{\mathbb{R}}^{n}$ be a proper convex cone, and suppose that $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$ is $\mathcal{K}$}-completely log-concave in the sense of Definition˜2.1. Then $f$ is linewise $\mathcal{K}$}-completely log-concave.
Proof.
Fix $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}\setminus\{0\}$, and let $I_{x,v}$ and $g_{x,v}$ be as in Definition˜3.12. We must show that $g_{x,v}$ is completely log-concave on $I_{x,v}$ in the univariate sense.
By Definition˜2.1, for every integer $m\geq 0$ the directional derivative
$D_{v}^{m}f\;=\;\underbrace{D_{v}\cdots D_{v}}_{m\ \text{times}}f$
is log-concave on $\operatorname{int}{\mathcal{K}}$, since each direction $v$ belongs to $\mathcal{K}$}. For $t\in I_{x,v}$ we have $x+tv\in\operatorname{int}{\mathcal{K}}$, and
$\frac{d^{m}}{dt^{m}}g_{x,v}(t)\;=\;\frac{d^{m}}{dt^{m}}f(x+tv)\;=\;D_{v}^{m}f(x+tv).$
Thus the $m$-th derivative of $g_{x,v}$ is the restriction of the log-concave function $D_{v}^{m}f$ (defined on $\operatorname{int}{\mathcal{K}}$) to the line segment $\{x+tv:t\in I_{x,v}\}$. Since restrictions of a log-concave function to a line are log-concave on their domain, it follows that $g_{x,v}^{(m)}$ is log-concave on $I_{x,v}$ for every $m\geq 0$.
This is precisely the definition of complete log-concavity for the univariate function $g_{x,v}$ on $I_{x,v}$, and therefore $f$ is linewise $\mathcal{K}$}-completely log-concave. ∎
Next we provide two necessary conditions for a $f$ to be $\mathcal{K}$}-CLC along $v\in\operatorname{int}{\mathcal{K}}$.
Proposition 3.14.
Let $f(x)$ be a nonzero $\mathcal{K}$}-CLC over a proper convex cone $\mathcal{K}$}. Then for any $x,v\in\operatorname{int}{\mathcal{K}}$, the coefficients of $f(x+tv)$ are positive.
Proof.
Since $f$ is CLC over $\mathcal{K}$}, its directional derivatives are log-concave, i.e., they satisfy the positivity condition at each degree level $d\geq 1$. Therefore, in particular, $D_{v}f(x+tv)=f^{{}^{\prime}}(x+tv)>0,$ and $D_{v}D_{v}f(x+tv)=f^{{}^{\prime\prime}}(x+tv)>0$ for all $v\in\operatorname{int}{\mathcal{K}}$. That enforces all the coefficients of $f(x+tv)=f(x)+tD_{v}f(x)+\frac{t^{2}}{2}D_{v}^{2}f(x)+\dots+\frac{t^{d}}{d!}D_{v}^{d}f(x)$ to be positive. ∎
Corollary 3.15.
Let $f(x)$ be a strictly $\mathcal{K}$}-CLC over a proper convex cone $\mathcal{K}$}. Then for any $x,v\in{\mathcal{K}}$, the coefficients of $f(x+tv)$ are positive. In particular, if $f(x)$ is a strictly $\mathcal{K}$}-Lorentzian polynomial over a proper convex cone $\mathcal{K}$}, then for any $x,v\in{\mathcal{K}}$, the coefficients of $f(x+tv)$ are positive.
Proof.
It follows from Definition˜2.1 and Proposition˜3.14. ∎
Theorem 3.16.
Let $K\subseteq{\mathbb{R}}^{n}$ be a proper convex cone and let $f(x)$ be a nonzero $K$-CLC polynomial. Let $d:=\deg f$ and let $x,v\in\operatorname{int}K$. Then the sequence
$\bigl\{f(x),\,D_{v}f(x),\,D_{v}^{2}f(x),\,\dots,\,D_{v}^{d}f(x)\bigr\}$
of positive real numbers is log-concave.
Proof.
Since $f$ is $K$-CLC in the sense of Definition˜2.1, it follows from Proposition˜3.13 that $f$ is linewise $K$-CLC in the sense of Definition˜3.12. Hence, for any $x,v\in\operatorname{int}K$, the univariate restriction
$g(t)\coloneqq f(x+tv)$
is completely log-concave on $t\geq 0$.
By Proposition 3.14, for any $x,v\in\operatorname{int}K$ all coefficients of $g(t)$ are positive. Writing
$g(t)=f(x+tv)=f(x)+tD_{v}f(x)+\frac{t^{2}}{2!}D_{v}^{2}f(x)+\cdots+\frac{t^{d}}{d!}D_{v}^{d}f(x),$
we see that the coefficients are
$a_{k}=\frac{1}{k!}D_{v}^{k}f(x),\qquad k=0,1,\dots,d.$
Since $g$ is univariate CLC with positive coefficients, Lemma 3.10 applies and implies that the sequence
$\{0!a_{0},\,1!a_{1},\,\dots,\,d!a_{d}\}$
is log-concave. Substituting $a_{k}=\frac{1}{k!}D_{v}^{k}f(x)$, we obtain
$0!a_{0}=f(x),\quad 1!a_{1}=D_{v}f(x),\quad\dots,\quad d!a_{d}=D_{v}^{d}f(x),$
so the sequence
$\bigl\{f(x),\,D_{v}f(x),\,D_{v}^{2}f(x),\,\dots,\,D_{v}^{d}f(x)\bigr\}$
is log-concave, as claimed. ∎
Remark 3.17.
The converse of Theorem˜3.16 does not hold. Even if, for every $x\in\operatorname{int}K$ and every $v\in\operatorname{int}K$, the sequence
$\bigl\{f(x),D_{v}f(x),\dots,D_{v}^{d}f(x)\bigr\}$
is log-concave, this does not imply that $f$ is $K$-CLC.
4. Half-Plane Property
Let $\mathcal{H}$} be the open left half-plane $\{x\in{\mathbb{C}}:\mathsf{Re}\ x<0\}$. A real valued $n\times n$ matrix $A$ is stable if all its eigenvalues belong to the open left half plane. A univariate polynomial $g(t)$ is called Hurwitz-stable if all its roots have negative real part.
Note that Hurwitz stability is closed under inversion, so we have
Proposition 4.1.
$g(t)\in{\mathbb{R}}[t]$ is Hurwitz-stable of degree $d$ over an interval $[a,b]$ if and only if $t^{d}g(1/t)$ is Hurwitz-stable over $[a,b]$.
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$. Then the corresponding $d\times d$ Hurwitz matrix is given by ${\mathrm{Hur}}_{f(t)}=\begin{bmatrix}a_{1}&a_{3}&a_{5}&\dots&\dots&\dots&0&0&0\\ a_{0}&a_{2}&a_{4}&&&&\vdots&\vdots&\vdots\\ 0&a_{1}&a_{3}&&&&\vdots&\vdots&\vdots\\ \vdots&a_{0}&a_{2}&\ddots&&&0&\vdots&\vdots\\ \vdots&0&a_{1}&&\ddots&&a_{d}&\vdots&\vdots\\ \vdots&\vdots&a_{0}&&&\ddots&a_{d-1}&0&\vdots\\ \vdots&\vdots&0&&&&a_{d-2}&a_{d}&\vdots\\ \vdots&\vdots&\vdots&&&&a_{d-3}&a_{d-1}&0\\ 0&0&0&\dots&\dots&\dots&a_{d-4}&a_{d-2}&a_{d}\end{bmatrix}$, cf., [stabilityHurwitz, Chap-7].
Theorem 4.2.
Using Routh-Hurwitz criterion, cf., [Bellmanselected, Chap 4], the FAE:
  1. (a)
    $f(t)$ is Hurwitz-stable
  2. (b)
    All the leading principal minors of the corresponding Hurwitz matrix ${\mathrm{Hur}}_{f(t)}$ are positive.
Remark 4.3.
By expanding the determinant of ${\mathrm{Hur}}_{f(t)}$ along the elements of its last column gives $\det({\mathrm{Hur}}_{f(t)})=a_{d}\Delta_{d-1}$ where $\Delta_{k}$ represents the $k\times k$ leading principal minor of ${\mathrm{Hur}}_{f(t)}$. Since $a_{i}>0$, to demonstrate $f(t)$ of degree $d$ is Hurwitz-stable, it suffices to verify that $\Delta_{k}$ are positive for all $1\leq k\leq d-1$.
The following result is a refinement of the Routh-Hurwitz criterion above.
Theorem 4.4.
(Liénard-Chipart) Necessary and sufficient conditions for the polynomial $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ with $a_{0}>0$ to be stable can be given in any one of the following four forms: [gantmacher, pg-221]
  1. (1)
    $a_{k}>0,a_{k-2}>0,\dots;\Delta_{1}>0,\Delta_{3}>0,\dots$
  2. (2)
    $a_{k}>0,a_{k-2}>0,\dots;\Delta_{2}>0,\Delta_{4}>0,\dots$
  3. (3)
    $a_{k}>0,a_{k-1}>0,a_{k-3}>0,\dots;\Delta_{1}>0,\Delta_{3}>0,\dots$
  4. (4)
    $a_{k}>0,a_{k-1}>0,a_{k-3}>0,\dots;\Delta_{2}>0,\Delta_{4}>0,\dots$
A polynomial $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$ is called (Hurwitz-)stable if $f(z)\neq 0$ for all $z=(z_{1},\dots,z_{n})\in{\mathbb{C}}^{n}$ with $\mathsf{Re}(z_{i})>0$ for all $1\leq i\leq n$. Equivalently, all zeros of $f$ lie in the closed left half-plane in each coordinate. By convention, we also regard the zero polynomial as stable.
Equivalently, if $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$ is homogeneous and nonzero, then $f$ is Hurwitz-stable if and only if, for all $x,v\in({\mathbb{R}}_{>0})^{n}$, the univariate polynomial
$t\ \longmapsto\ f(x+tv)$
is a Hurwitz-stable polynomial in $t$, that is, all of its zeros have negative real part; see, e.g., [Gregor, Lemma 6], [Oxley, Prop. 5.2].
Definition 4.5.
Let ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a proper convex cone with nonempty interior $\operatorname{int}{\mathcal{K}}$. A polynomial $f\in{\mathbb{R}}_{n}^{d}[x]$ is said to be Hurwitz-stable over $\mathcal{K}$} if either $f\equiv 0$, or else for every $x,v\in\operatorname{int}{\mathcal{K}}$ the univariate polynomial
$t\ \longmapsto\ f(x+tv)$
is Hurwitz-stable in the univariate sense, i.e., all of its zeros have negative real part.
Theorem 4.6.
[Gulerhyperbolicconvex, Theorem 5.3] Any hyperbolic polynomial $f\in{\mathbb{R}}[x]^{d}_{n}$ is Hurwitz-stable over its hyperbolicity cone $\Lambda_{++}(f,e)$ containing $e$ such that $f(e)>0$.
Here we derive Newton-type inequalities for directional derivatives, which we then feed into stability criteria.
Proposition 4.7.
Let $f(t)=\sum_{k=0}^{n}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,n$. If $a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1}$ for all $1\leq k\leq d-1$, the following inequalities hold:
  1. (a)
    For any $1\leq k\leq d-2$, $a_{k}a_{k+1}\geq\frac{k+2}{k}a_{k-1}a_{k+2}\Leftrightarrow\frac{a_{k}}{a_{k-1}}\geq\frac{k+2}{k}\frac{a_{k+2}}{a_{k+1}}\Leftrightarrow\frac{a_{k+1}}{a_{k-1}}\geq\frac{k+2}{k}\frac{a_{k+2}}{a_{k}}$.
  2. (b)
    For any $1\leq k\leq d-2$, $a_{k}^{2}\geq\frac{(k+2)(k+1)}{k(k-1)}a_{k-2}a_{k+2}$.
  3. (c)
    For any $1\leq k\leq d-3$, $a_{k}a_{k+2}\geq\frac{k+3}{k}a_{k-1}a_{k+3}$, and for any $1\leq k\leq d-1$, $a_{k}a_{d-1}\geq\frac{d}{k}a_{k-1}a_{d}$.
  4. (d)
    $a_{i}a_{j}\geq\frac{l}{i}a_{k}a_{l}$ if $i+j=k+l$ for any $k<i<j<l$.
Proof.
$(a)$
$a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1},a_{k+1}^{2}\geq\frac{k+2}{k+1}a_{k}a_{k+2}\Rightarrow a_{k}a_{k+1}\geq\frac{k+2}{k}a_{k-1}a_{k+2}$
$(b)$
$a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1},a_{k+1}^{2}\geq\frac{k+2}{k+1}a_{k}a_{k+2},a_{k-1}^{2}\geq\frac{k}{k-1}a_{k-2}a_{k}$
$\Rightarrow a_{k}^{2}\geq\frac{k+1}{k}\sqrt{\frac{k+2}{k+1}}\sqrt{\frac{k}{k-1}}\sqrt{a_{k-2}a_{k}}\sqrt{a_{k}a_{k+2}}$
$\Rightarrow a_{k}^{2}\geq\frac{k+1}{k}\frac{k+2}{k-1}a_{k-2}a_{k+2}$
$(c)$
$a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1},a_{k+1}^{2}\geq\frac{k+2}{k+1}a_{k}a_{k+2},a_{k+2}^{2}\geq\frac{k+3}{k+2}a_{k+1}a_{k+3}$
$\Rightarrow a_{k}^{2}a_{k+2}^{2}\geq\frac{k+1}{k}\frac{k+3}{k+2}a_{k-1}a_{k+1}^{2}a_{k+3}$
$\Rightarrow a_{k}^{2}a_{k+2}^{2}\geq\frac{k+1}{k}\frac{k+3}{k+2}\frac{k+2}{k+1}a_{k-1}a_{k}a_{k+2}a_{k+3}$
$\Rightarrow a_{k}a_{k+2}\geq\frac{k+3}{k}a_{k-1}a_{k+3}$
By applying the same process iteratively, we arrive at the final inequality of $(c)$.
$(d)$ This result follows from the established pattern above. ∎
Proposition 4.8.
Let $f(t)=\sum_{k=0}^{n}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,n$. If $a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1}$ for all $1\leq k\leq d-1$, then the following inequalities hold:
  1. (a)
    Monotonic chains of adjacent ratios: $\frac{a_{2}}{a_{0}}>\frac{a_{3}}{a_{1}}>\frac{a_{4}}{a_{2}}>\frac{a_{5}}{a_{3}}>\dots>\frac{a_{k}}{a_{k-2}}>\frac{a_{k+1}}{a_{k-1}}>\frac{a_{k+2}}{a_{k}}>\dots$.
  2. (b)
    Separated-index ratio chains: $\frac{a_{1}}{a_{0}}>\frac{a_{3}}{a_{2}}>\frac{a_{5}}{a_{4}}>\dots\frac{a_{k-1}}{a_{k-2}}>\frac{a_{k+1}}{a_{k}}>\frac{a_{k+3}}{a_{k+2}}>\dots$
  3. (c)
    Cross-ratio type inequalities: $\frac{a_{k+1}-a_{k}r}{a_{k-1}-a_{k-2}r}<\frac{a_{k+1}}{a_{k-1}}\Leftrightarrow a_{k}a_{k-1}>a_{k-2}a_{k+1}$, and $\frac{a_{k+1}-a_{k+2}r}{a_{k-1}-a_{k}r}>\frac{a_{k+1}}{a_{k-1}}\Leftrightarrow a_{k}a_{k+1}>a_{k-1}a_{k+2}$ for any positive $r$.
Proof.
$(a)$ The chain of inequalities is obtained by combining Proposition˜4.7 $(a)$ and $(d)$.
$(b)$ It follows from $(a)$ by rearranging the position of $a_{i}$.
$(c)$ These are derived by applying the identity $\frac{a}{b}>\frac{c}{d}\Leftrightarrow ad>bc$.
Note that strong log-concavity $\Rightarrow$ very rigid ordering for coefficient ratios and related rational expressions. We translate abstract log-concavity into very concrete, ordered ratio inequalities that can be plugged directly into determinant conditions for Hurwitz matrices.
Lemma 4.9.
Let $f(x)$ be a nonzero $\mathcal{K}$}-CLC over a proper convex cone $\mathcal{K}$} and $v\in\operatorname{int}{\mathcal{K}}$. Then the following inequalities are satisfied.
  1. (a)
    $(D_{v}^{k}f(x))^{2}\geq D_{v}^{k-1}f(x)D_{v}^{k+1}f(x)$ for any $1\leq k\leq d-1$ and $x\in\operatorname{int}{\mathcal{K}}$.
  2. (b)
    For any $0\leq k\leq d-3$, $D_{v}^{k+1}f(x)D_{v}^{k+2}f(x)\geq D_{v}^{k}f(x)D_{v}^{k+3}f(x)\Leftrightarrow\frac{D_{v}^{k+1}f(x)}{D_{v}^{k}f(x)}\geq\frac{D_{v}^{k+3}f(x)}{D_{v}^{k+2}f(x)}$.
  3. (c)
    For any $0\leq k\leq d-4$, $(D_{v}^{k+2}f(x))^{2}\geq D_{v}^{k}f(x)D_{v}^{k+4}f(x)$.
Proof.
$(a)$ Since $f$ is $\mathcal{K}$}-CLC, so the sequence $\{f(x),D_{v}f(x),D_{v}^{2}f(x),\dots,D_{v}^{d}f(x)\}$ of positive real numbers forms a log-concave sequence for any $x\in\operatorname{int}{\mathcal{K}}$ by Theorem˜3.16.
$(b)$ and $(c)$ The inequalities hold for any $x\in\operatorname{int}{\mathcal{K}}$ by Proposition˜4.7. ∎
Recall that two polynomials $f$ and $g$ are Hurwitz-stable if and only if the product $fg$ is Hurwitz-stable. Additionally, the Hadamard (coefficient-wise) product of two Hurwitz stable polynomials is again Hurwitz stable. Here we report the product rule for $\mathcal{K}$}-CLC polynomials.
Proposition 4.10.
[Hereditary, Corollary 8:13] If $f\in{\mathbb{R}}[x]$ and $g\in{\mathbb{R}}[x]$ are $\mathcal{K}$}-CLC multivariate polynomials, the product $fg$ is $\mathcal{K}$}-CLC polynomial
The converse of the above statement need not be true in general.
Example 4.11.
Although $f(t)=3+7t+7t^{2}+4t^{3}=(1+t+t^{2})(3+4t)$ is CLC on $t\geq 0$ but its factor $1+t+t^{2}$ is not CLC over $t\geq 0$.
We establish a connection between $\mathcal{K}$}-CLC polynomials and Hurwitz-stable polynomials over $\mathcal{K}$}. Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ be a real univariate polynomial of degree $d$. The even and odd parts of a real polynomial $f(t)$ are defined as:
$f^{\mathrm{even}}(t)$ $:=a_{0}+a_{2}t^{2}+a_{4}t^{4}+\dots$
$f^{\mathrm{odd}}(t)$ $:=a_{1}t+a_{3}t^{3}+a_{5}t^{5}+\dots$
Define
$f^{e}(t):=\sum_{m=0}^{\lfloor{d/2\rfloor}}(-1)^{m}a_{2m}t^{m}$
$f^{o}(t):=\sum_{m=0}^{\lfloor{(d-1)/2\rfloor}}(-1)^{m}a_{2m+1}t^{m}$
Theorem 4.12.
Hermite-Biehler’s criterion for stability: [Levindistribution, Chap VII], [Kurtzsufficient] Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,d$. $f(t)$ is stable if and only if $f^{e}(t)$ and $tf^{o}(t)$ have simple real interlacing zeros. Alternatively, $f(t)$ is stable if and only if all the zeros of $f^{\mathrm{even}}(t)$ and $f^{\mathrm{odd}}(t)$ are distinct, lie on the imaginary axis and alternate along it.
We use the following Hurwitz-Stability test for real polynomials based on the Interlacing Theorem and therefore, on the Boundary Crossing Theorem, cf. [Bhattacharyyarobust, Chap 1]. Note that $f(t)=f^{\mathrm{even}}(t)+f^{\mathrm{odd}}(t)$. Define the polynomial $g(t)$ of degree d1d-1 by :
$\mbox{If}\ d=2m:g(t)$ $=[f^{\mathrm{even}}(t)-\frac{a_{2m}}{a_{2m-1}}tf^{\mathrm{odd}}(t)]+f^{\mathrm{odd}}(t)]$
$\mbox{If}\ d=2m+1:g(t)$ $=[f^{\mathrm{odd}}(t)-\frac{a_{2m+1}}{a_{2m}}tf^{\mathrm{even}}(t)]+f^{\mathrm{even}}(t)]$
i.e., in general with $\mu=\frac{a_{d}}{a_{d-1}}$,
$g(t)=a_{d-1}t^{d-1}+(a_{d-2}-\mu a_{d-3})t^{d-2}+a_{d-3}t^{d-3}+(a_{d-4}-\mu a_{d-5})t^{d-4}+\dots$
Theorem 4.13.
[Bhattacharyyarobust, Chap 1] If $f(t)$ has all its coefficients positive, $f(t)$ is stable if and only if $g(t)$ is stable
It is well known and easy to verify that polynomials of degree $1$ and $2$ with positive leading coefficients are Hurwitz-stable if and only if all their coefficients are positive. Thus the result is trivially true for $d=1$ and $d=2$. By Theorem˜4.2, it’s easy to verify that for $d=3$, the $3\times 3$ leading principal minor of ${\mathrm{Hur}}_{f(t)}$ is positive if and only if $a_{1}a_{2}-a_{0}a_{3}>0$, cf. [Bhattacharyyarobust, Chap-1]. Since $a_{1}a_{2}>3a_{0}a_{3}$ by Proposition˜4.7, so $f(t)$ is Hurwitz-stable.
Proposition 4.14.
Let $f(t)=\sum_{k=0}^{4}a_{k}t^{k}$ be a quartic polynomial with $a_{k}>0$ for $k=0,\dots,4$. Suppose that
$a_{k}^{2}\;>\;\frac{k+1}{k}\,a_{k-1}a_{k+1}\qquad\text{for }k=1,2,3.$
Then $f(t)$ is Hurwitz–stable.
Proof.
By Theorem˜4.13, $f(t)=\sum_{k=0}^{4}a_{k}t^{k}$ is Hurwitz–stable if and only if
$g(t)\;:=\;f(t)-\mu tf^{\mathrm{odd}}(t)\;=\;a_{0}+a_{1}t+\Bigl(a_{2}-\frac{a_{4}}{a_{3}}a_{1}\Bigr)t^{2}+a_{3}t^{3}$
is Hurwitz–stable, where $\mu=a_{4}/a_{3}$. Set
$b_{0}:=a_{0},\quad b_{1}:=a_{1},\quad b_{2}:=a_{2}-\frac{a_{4}}{a_{3}}a_{1},\quad b_{3}:=a_{3},$
so $g(t)=b_{0}+b_{1}t+b_{2}t^{2}+b_{3}t^{3}$.
First, $b_{2}>0$ by Proposition˜4.7(a), which gives
$a_{2}a_{3}\;>\;a_{1}a_{4}\quad\Longrightarrow\quad a_{2}-\frac{a_{4}}{a_{3}}a_{1}>0.$
Hence $b_{0},b_{1},b_{2},b_{3}$ are all positive.
For a cubic with positive coefficients $b_{0},b_{1},b_{2},b_{3}>0$, the Routh–Hurwitz criterion says that Hurwitz stability is equivalent to
$b_{1}b_{2}\;>\;b_{0}b_{3}\quad\Longleftrightarrow\quad a_{1}\Bigl(a_{2}-\frac{a_{4}}{a_{3}}a_{1}\Bigr)\;>\;a_{0}a_{3}.$
Equivalently,
$a_{1}a_{2}a_{3}-a_{1}^{2}a_{4}-a_{0}a_{3}^{2}\;>\;0.$
Thus it remains to show
(1) $a_{1}a_{2}a_{3}\;>\;a_{1}^{2}a_{4}+a_{0}a_{3}^{2}.$
Define the dimensionless quantities
$r_{1}:=\frac{a_{1}^{2}a_{4}}{a_{1}a_{2}a_{3}}=\frac{a_{1}a_{4}}{a_{2}a_{3}},\qquad r_{2}:=\frac{a_{0}a_{3}^{2}}{a_{1}a_{2}a_{3}}=\frac{a_{0}a_{3}}{a_{1}a_{2}}.$
Then (1) is equivalent to
$1-r_{1}-r_{2}\;>\;0.$
From the assumed inequalities
$a_{1}^{2}>2a_{0}a_{2},\qquad a_{2}^{2}>\frac{3}{2}a_{1}a_{3},\qquad a_{3}^{2}>\frac{4}{3}a_{2}a_{4},$
we obtain the following bounds.
1. For $r_{2}$:
$\frac{a_{0}}{a_{1}}<\frac{a_{1}}{2a_{2}}\quad\Rightarrow\quad r_{2}=\frac{a_{0}}{a_{1}}\cdot\frac{a_{3}}{a_{2}}<\frac{a_{1}}{2a_{2}}\cdot\frac{a_{3}}{a_{2}}=\frac{a_{1}a_{3}}{2a_{2}^{2}}.$
Using $a_{2}^{2}>\frac{3}{2}a_{1}a_{3}$ gives
$\frac{a_{1}a_{3}}{a_{2}^{2}}<\frac{2}{3}\quad\Rightarrow\quad r_{2}<\frac{1}{2}\cdot\frac{2}{3}=\frac{1}{3}.$
2. For $r_{1}$:
$r_{1}=\frac{a_{1}}{a_{2}}\cdot\frac{a_{4}}{a_{3}}.$
From $a_{2}^{2}>\frac{3}{2}a_{1}a_{3}$, $\frac{a_{1}}{a_{2}}<\frac{2}{3}\cdot\frac{a_{2}}{a_{3}}$, and from $a_{3}^{2}>\frac{4}{3}a_{2}a_{4}$, $\frac{a_{4}}{a_{3}}<\frac{3}{4}\cdot\frac{a_{3}}{a_{2}}$. Multiplying these estimates gives
$r_{1}<\frac{2}{3}\cdot\frac{3}{4}\cdot\frac{a_{2}}{a_{3}}\cdot\frac{a_{3}}{a_{2}}=\frac{1}{2}.$
Combining the two bounds, we get
$r_{1}+r_{2}<\frac{1}{2}+\frac{1}{3}=\frac{5}{6}<1,$
so $1-r_{1}-r_{2}>0$, which is exactly (1). Hence $g$ is Hurwitz–stable, and therefore $f$ is Hurwitz–stable by Theorem˜4.13. ∎
Corollary 4.15.
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ be a univariate strictly CLC polynomial for $t\geq 0$ with $a_{k}>0$ for $k=0,\dots,d$. If $1\leq d\leq 4$, then $f(t)$ is Hurwitz–stable.
Lemma 4.16.
Assume $a_{0},a_{1},a_{2},a_{3},a_{4},a_{5}>0$ and $a_{2}-a_{4}k>0$. For any positive $k$,
$\frac{a_{1}}{a_{0}}>\frac{a_{3}-a_{5}k}{a_{2}-a_{4}k}\quad\Longleftrightarrow\quad k<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}},$
provided $a_{1}a_{4}-a_{0}a_{5}>0$.
Proof.
By the standard identity for positive denominators,
$\frac{a}{b}>\frac{c}{d}\;\Longleftrightarrow\;ad>bc.$
Applying this with $a=a_{1}$, $b=a_{0}$, $c=a_{3}-a_{5}k$, $d=a_{2}-a_{4}k$ and rearranging gives
$a_{1}(a_{2}-a_{4}k)>a_{0}(a_{3}-a_{5}k)\;\Longleftrightarrow\;(a_{1}a_{4}-a_{0}a_{5})k<a_{1}a_{2}-a_{0}a_{3},$
which is equivalent to the claimed inequality when $a_{1}a_{4}-a_{0}a_{5}>0$. ∎
Theorem 4.17.
Let $f(t)=\sum_{k=0}^{5}a_{k}t^{k}$ be a univariate polynomial with $a_{k}>0$ for $k=0,\dots,5$. Suppose that
$a_{k}^{2}>\frac{k+1}{k}\,a_{k-1}a_{k+1}\quad\text{for }k=1,\dots,4,$
and
$(a_{1}a_{4}-a_{0}a_{5})^{2}<(a_{1}a_{2}-a_{0}a_{3})(a_{3}a_{4}-a_{2}a_{5}).$
Then $f(t)$ is Hurwitz-stable.
Proof.
By Theorem˜4.2 and Remark˜4.3, for $d=5$ it is sufficient to show that all leading principal minors $\Delta_{k}>0,\quad k=1,\dots,4,$ of the Hurwitz matrix
${\mathrm{Hur}}_{f(t)}=\begin{bmatrix}a_{1}&a_{3}&a_{5}&0&0\\ a_{0}&a_{2}&a_{4}&0&0\\ 0&a_{1}&a_{3}&a_{5}&0\\ 0&a_{0}&a_{2}&a_{4}&0\\ 0&0&a_{1}&a_{3}&a_{5}\end{bmatrix}$.
The Newton-type inequalities $a_{k}^{2}>\frac{k+1}{k}\,a_{k-1}a_{k+1},\quad k=1,\dots,4,$ imply in particular that
$a_{1}>0,\quad a_{1}a_{2}-a_{0}a_{3}>0,\quad\det\begin{bmatrix}a_{1}&a_{3}&0\\ a_{0}&a_{2}&a_{4}\\ 0&a_{1}&a_{3}\end{bmatrix}>0.$
From this it follows that $\det\begin{bmatrix}a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\\ 0&a_{1}&a_{3}\end{bmatrix}>0$, so $\Delta_{k}>0$ for $k=1,2,3$.
It remains to show that the $4\times 4$ leading principal minor is positive, i.e.,
(2) $\frac{a_{1}}{a_{0}}>\frac{\det\begin{bmatrix}a_{3}&a_{5}&0\\ a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\end{bmatrix}}{\det\begin{bmatrix}a_{2}&a_{4}&0\\ a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\end{bmatrix}}=\frac{a_{3}-a_{5}r}{a_{2}-a_{4}r},$
where $r:=\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}$. By the Newton-type inequalities and Proposition˜4.8(a),(c), we have
$\frac{a_{3}-a_{5}r}{a_{2}-a_{4}r}>\frac{a_{3}}{a_{2}}.$
On the other hand, by Lemma˜4.16, the inequality (2) holds if and only if
$r<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}},$
that is,
$\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}}\;\Longleftrightarrow\;(a_{1}a_{4}-a_{0}a_{5})^{2}<(a_{1}a_{2}-a_{0}a_{3})(a_{3}a_{4}-a_{2}a_{5}),$
which is exactly our additional hypothesis. Hence the $4\times 4$ leading principal minor is positive and therefore $f(t)$ is Hurwitz-stable. ∎
In fact, a stronger result is presented in [Hurwitzsuff, Theorem 3] for entire functions. However, we limit our focus here to univariate polynomials, cf. [Hurwitzsuff, Theorem 1].
Theorem 4.18.
If the coefficients of $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ are positive and satisfy
$a_{k}a_{k+1}\;\geq\;\alpha\,a_{k-1}a_{k+2},\qquad 1\leq k\leq d-1,$
where $\alpha$ is the unique positive root of the cubic
$t^{3}-t^{2}-2t-1=0\quad(\alpha\approx 2.1479),$
then $f(t)$ is Hurwitz-stable. In particular, the conclusion holds if
$a_{k}^{2}\;\geq\;\sqrt{\alpha}\,a_{k-1}a_{k+1},\qquad 1\leq k\leq d-1.$
Now we extend the above univariate results to multivariate homogeneous polynomials.
Theorem 4.19.
Let $f(x)\in{\mathbb{R}}^{d}_{n}[x]$ be a strictly $\mathcal{K}$}-Lorentzian polynomial of degree $d$ with $1\leq d\leq 4$ over a proper convex cone $\mathcal{K}$}. Then $f(x)$ is Hurwitz-stable over $\mathcal{K}$}.
Proof.
Since $f$ is strictly $\mathcal{K}$}-Lorentzian, Definition˜3.12 together with Theorem˜2.9 (the equivalence of strictly $\mathcal{K}$}-CLC and strictly $\mathcal{K}$}-Lorentzian polynomials in the homogeneous setting) implies that for any $x,v\in\operatorname{int}{\mathcal{K}}$ the univariate restriction
$f(x+tv)=\sum_{k=0}^{d}\frac{D_{v}^{k}f(x)}{k!}\,t^{k}$
is a strictly Lorentzian polynomial in $t$ for all $t$ such that $x+tv\in{\mathcal{K}}$. Moreover, by Corollary˜3.15, all coefficients $a_{k}:=\frac{D_{v}^{k}f(x)}{k!}$ are positive for $x,v\in\operatorname{int}{\mathcal{K}}$.
Since $f$ is strictly $\mathcal{K}$}-Lorentzian, Theorem˜3.16 yields the Newton-type inequalities
$\bigl(D_{v}^{k}f(x)\bigr)^{2}>\frac{k+1}{k}\,D_{v}^{k-1}f(x)\,D_{v}^{k+1}f(x),\qquad 1\leq k\leq d-1,$
for all $x,v\in\operatorname{int}{\mathcal{K}}$. In terms of the coefficients $a_{k}=\frac{D_{v}^{k}f(x)}{k!}$, this is exactly
$a_{k}^{2}>\frac{k+1}{k}\,a_{k-1}a_{k+1},\qquad 1\leq k\leq d-1.$
Therefore, by Proposition˜4.14, the univariate polynomial $t\mapsto f(x+tv)$ is Hurwitz-stable for every $x,v\in\operatorname{int}{\mathcal{K}}$. By Definition˜4.5, this means that $f(x)$ is Hurwitz-stable over $\mathcal{K}$}. ∎
Theorem 4.20.
Let $f(x)\in{\mathbb{R}}^{d}_{n}[x]$ be strictly $\mathcal{K}$}-Lorentzian of degree $d\geq 5$ over a proper convex cone $\mathcal{K}$}. Suppose that
$\bigl(D_{v}^{k}f(x)\bigr)^{2}>\frac{\sqrt{\alpha}\,k}{k+1}\,D_{v}^{k-1}f(x)\,D_{v}^{k+1}f(x)\quad\text{for all }1\leq k\leq d-1,\;x,v\in\operatorname{int}{\mathcal{K}},$
where $\alpha$ is the unique positive root of the cubic $t^{3}-t^{2}-2t-1=0$ (so $\alpha\approx 2.1479$). Then $f(x)$ is Hurwitz-stable over $\mathcal{K}$}.
Proof.
Fix $x,v\in\operatorname{int}{\mathcal{K}}$ and consider the univariate restriction
$g(t):=f(x+tv)=\sum_{k=0}^{d}a_{k}t^{k},\qquad\text{with }a_{k}:=\frac{D_{v}^{k}f(x)}{k!}.$
By Definition˜3.12 and Theorem˜2.9, $g$ is strictly Lorentzian, so in particular $a_{k}>0$ for all $k$.
The assumed inequalities on $D_{v}^{k}f(x)$ translate into the coefficient inequalities required in Theorem˜4.18. Indeed,
$a_{k}^{2}=\frac{\bigl(D_{v}^{k}f(x)\bigr)^{2}}{(k!)^{2}}>\sqrt{\alpha}\,\frac{k}{k+1}\frac{D_{v}^{k-1}f(x)\,D_{v}^{k+1}f(x)}{(k!)^{2}}=\sqrt{\alpha}\,\frac{D_{v}^{k-1}f(x)}{(k-1)!}\,\frac{D_{v}^{k+1}f(x)}{(k+1)!}=\sqrt{\alpha}\,a_{k-1}a_{k+1}.$
Thus the sufficient condition in Theorem˜4.18 is satisfied, and so $g(t)$ is Hurwitz-stable. Since this holds for every $x,v\in\operatorname{int}{\mathcal{K}}$, it follows from Definition˜4.5 that $f(x)$ is Hurwitz-stable over $\mathcal{K}$}. ∎
Remark 4.21.
Not every strictly $\mathcal{K}$}-CLC polynomial is Hurwitz-stable over $\mathcal{K}$}. For instance, for ${\mathcal{K}}={\mathbb{R}}_{\geq 0}$, the univariate polynomial
$f(t)=5+14t+12.5t^{2}+7.2t^{3}+3t^{4}+t^{5}$
is Lorentzian for $t\geq 0$ but not Hurwitz-stable. Conversely, the polynomial
$f(t)=5+25t+50t^{2}+30t^{3}+10t^{4}+3t^{5}$
is Hurwitz-stable but not Lorentzian. These examples show that the sufficient conditions in Theorems˜4.19 and 4.20 are not necessary.
5. Applications: Stability Analysis for LTI Systems
Consider the linear time-invariant (LTI) system
$\dot{x}=Ax,\qquad A\in{\mathbb{R}}^{n\times n}.$
A matrix $A$ is called stable (or Hurwitz) if every eigenvalue of $A$ has negative real part, i.e., all eigenvalues lie in the open left half-plane. If $A$ is real symmetric, then $A$ is stable if and only if $A$ is negative definite. Lyapunov’s theorem asserts that $A$ is stable if and only if there exists a symmetric positive definite matrix $P$ such that the Lyapunov operator
$L_{A}(P):=A^{\top}P+PA$
is negative definite.
Theorem 5.1 (Routh–Hurwitz criterion).
Let $A\in{\mathbb{R}}^{n\times n}$, and let $\chi_{A}(t)$ be its characteristic polynomial. Then the following are equivalent [Poznyak, Chap. 9], [stabilityHurwitz, Chap. 7]:
  1. (a)
    The matrix $A$ is stable, i.e., all eigenvalues of $A$ lie in the open left half-plane.
  2. (b)
    The characteristic polynomial $\chi_{A}(t)$ is Hurwitz-stable.
  3. (c)
    All the leading principal minors of the Hurwitz matrix ${\mathrm{Hur}}_{\chi_{A}(t)}$ are positive.
Hence we have the following.
Theorem 5.2.
Let $f(t)$ be a strictly CLC polynomial of degree $d\leq 4$ for $t\geq 0$. Then the corresponding dynamical system $\frac{dx}{dt}=Ax$ is stable if the characteristic polynomial $\chi(A)=f(t)$.
Proof.
Since $f(t)$ is a strictly CLC polynomial of degree $d\leq 4$ for $t\geq 0$, $f(t)$ is Hurwitz-stable by Theorem˜4.19. Then it follows from Theorem˜5.1. ∎
Corollary 5.3.
Let $f(t)=\sum_{k=0}^{5}a_{k}t^{k}$ be a strictly CLC polynomial of degree $5$ for $t\geq 0$. If
$(a_{1}a_{4}-a_{0}a_{5})^{2}<(a_{1}a_{2}-a_{0}a_{3})\,(a_{3}a_{4}-a_{2}a_{5}),$
and $\chi_{A}(t)=f(t)$ is the characteristic polynomial of a matrix $A$, then the LTI system $\dot{x}=Ax$ is stable.
Proof.
By Theorem 4.17, the inequality implies that $f(t)$ is Hurwitz-stable. Hence $\chi_{A}(t)$ is Hurwitz-stable, and the claim follows from Theorem 5.1. ∎
Theorem 5.4.
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k},d\geq 5$ be a univariate polynomial with positive coefficients and $a_{k}^{2}>\sqrt{\alpha}a_{k-1}a_{k+1}$ where $\alpha$ is the unique positive root of the polynomial $t^{3}-t^{2}-2t-1(\alpha\approx 2.1479)$. Then the corresponding dynamical system $\frac{dx}{dt}=Ax$ is stable if the characteristic polynomial $\chi(A)=f(t)$,.
Proof.
It follows directly from Theorem 4.18 that $f(t)$ is Hurwitz-stable since the assumption $a_{k}^{2}>\sqrt{\alpha}\,a_{k-1}a_{k+1}$ for all $1\leq k\leq d-1$ implies the sufficient condition in that theorem. Then it follows from Theorem 5.1. ∎
Furthermore, we extend the result for multivariate polynomials.
Theorem 5.5.
Let $f(x)$ be a strictly $\mathcal{K}$}-Lorentzian polynomial of degree $d\leq 4$ over a proper convex cone $\mathcal{K}$}. Suppose there exist $x_{0},v\in\operatorname{int}{\mathcal{K}}$ such that the characteristic polynomial of a matrix $A$ is given by
$\chi_{A}(t)=f(x_{0}+tv).$
Then the LTI system $\dot{x}=Ax$ is stable.
Proof.
By Theorem 4.19, $f$ is Hurwitz-stable over $\mathcal{K}$}, so for any $x,v\in\operatorname{int}{\mathcal{K}}$ the univariate restriction $t\mapsto f(x+tv)$ is Hurwitz-stable. In particular, $\chi_{A}(t)=f(x_{0}+tv)$ is Hurwitz-stable. The conclusion then follows from Theorem 5.1. ∎
Theorem 5.6.
Let $f(x)$ be a strictly $\mathcal{K}$}-Lorentzian polynomial of degree $d\geq 5$ over a proper convex cone $\mathcal{K}$}. Suppose there exist $x_{0},v\in\operatorname{int}{\mathcal{K}}$ such that
$\chi_{A}(t)=f(x_{0}+tv)=\sum_{k=0}^{d}a_{k}t^{k}$
is the characteristic polynomial of a matrix $A$, and the coefficients satisfy
$a_{k}^{2}>\sqrt{\alpha}\,a_{k-1}a_{k+1},\qquad 1\leq k\leq d-1,$
where $\alpha$ is the unique positive root of $t^{3}-t^{2}-2t-1=0$ ($\alpha\approx 2.1479$). Then the LTI system $\dot{x}=Ax$ is stable.
Proof.
By the coefficient inequalities and Theorem 4.18, $\chi_{A}(t)$ is Hurwitz-stable. Stability of $\dot{x}=Ax$ then follows from Theorem 5.1. ∎
Remark 5.7.
Whenever the characteristic polynomial $\chi_{A}(t)$ of a stable matrix $A$ can be realized as a univariate restriction $\chi_{A}(t)=f(x_{0}+tv)$ of a Hurwitz-stable polynomial $f$ over a proper convex cone $\mathcal{K}$} with $x_{0},v\in\operatorname{int}{\mathcal{K}}$, we may regard $f$ as a $\mathcal{K}$}-Lorentzian lift of the dynamical system $\dot{x}=Ax$. This viewpoint connects LTI stability with the cone-geometric properties developed in this paper.
In particular, for a linear time-invariant system $\dot{x}=Ax$, Hurwitz stability of the characteristic polynomial $\chi_{A}(t)$ is equivalent to asymptotic stability of the equilibrium $x=0$. Thus, our Newton-type coefficient inequalities for strictly CLC and strictly $\mathcal{K}$}-Lorentzian polynomials yield simple sufficient tests ensuring stability of LTI systems directly from the coefficients of $\chi_{A}(t)$. Conversely, given a strictly $\mathcal{K}$}-CLC polynomial with suitable Hurwitz-stability properties, we can construct LTI systems whose spectral behaviour is encoded by $f$.
Acknowledgment: During the preparation of this manuscript, the author used ChatGPT to improve the clarity and conciseness of the writing. All content was subsequently reviewed, revised, and verified by the author, who takes full responsibility for the final version of the article.